# Step-Controlled DPO: Leveraging Stepwise Error for Enhanced Mathematical Reasoning

### v1

@@ 1 Introduction
Large language models (LLMs) have shown great potential in mathematical problem-solving. Recently, Direct Preference Optimization (DPO; (26)) has emerged as a popular choice for aligning LLMs with relative feedback to improve the quality of generated text. Prior works (8; 24; 41) have demonstrated that reinforcement learning algorithms and DPO can improve the mathematical reasoning abilities of LLMs, making the generated reasoning process more controllable. Different from other tasks that need human or AI feedback, the final answer to a mathematical problem serves as a reliable way to judge the quality of the model’s response, since a mathematical problem typically has a single correct answer. As a result, the responses producing the correct final answers are desirable and can serve as the preferred samples, while the ones reaching incorrect final answers are undesirable and can serve as the dispreferred samples.
@@ 2 Related Work
Aligning LLMs Using Relative Feedback. Methods that align LLMs with human or AI annotated preference data have been used to improve performance on a variety of downstream tasks such as translation (15), summarization (31; 54), and instruction-following (23; 27). Reinforcement learning from human (or AI) feedback (8; 4) first trains a reward model, then uses reinforcement learning algorithms such as REINFORCE (40), PPO (28), or variants (27) to maximize the reward. To simplify the pipeline, several direct alignment methods (26; 1; 52) have been proposed. Among them, DPO (26) and several of its variants (24; 10; 17; 1) offer a way to optimize the reward function without having to train an extra reward model, proving highly effective on various tasks (33; 48).
Recently, these preference alignment methods have also been applied to mathematical problem-solving tasks. DeepSeekMath (29) uses RL to improve mathematical accuracy, while ChatGLM-Math (41) uses DPO to improve model’s mathematical generation quality. Process supervision (16) uses stepwise preference of mathematical solution in its RL finetuning, which is highly effective but needs costly fine-grained human annotation. Our work offers a way to create stepwise error annotations of preferred and dispreferred solution pairs, and uses the data to improve DPO’s performance on mathematical problem-solving tasks.
@@ 3 Step-Controlled DPO Pipeline
We experiment with two solution formats: code-integrated solution format (53), and chain-of-thought solution format (38). For the code-integrated solution format, $t_{i}$ in $(t_{0},\ldots,t_{m})$ alternates between a code reasoning step and a natural language reasoning step, while for the chain-of-thought solution format, all the steps are in natural language. We primarily use the code-integrated solution format, as previous works (35; 19; 12) show that it results in higher accuracy than chain-of-thought. We finetune a Mistral-7B model with 34K samples of code-integrated solutions from GSM8K and 47K from MATH to create the initial model. We also validate our method on the chain-of-thought format using the off-the-shell MetaMath-Mistral-7B model, as well as MathCoder-Mistral-7B, which we trained using the MathCodeInstruct dataset (35).
@@ 5.2 Scaling of Data Amount on 20B Model
Training Settings. We use InternLM2-20B (6) as the foundation model, as it has demonstrated high performance in previous works (19; 6), even surpassing larger models such as Mixtral-8x7B (14) and Llama2-70B (32) in some cases. In the SFT stage, we finetune the model with a learning rate of $1.0\times 10^{-5}$ for 3 epochs, with a context length of 2048 tokens. In DPO and SCDPO training, we use a learning rate of $1.5\times 10^{-7}$ to train the SFT model for 2 epochs, with a context length of 1024 and $\beta$ set to 0.1. The models are trained on 16 NVIDIA A800 80GB GPUs with a batch size of 64.
@@ 6 Limitations and Future Work
Our work contains the following limitations, and we leave them for future work. Firstly, our work is conducted on purely linguistic models, which struggle to solve mathematical problems requiring an understanding of images. For example, questions in the geometry subject of the MATH dataset exhibit lower accuracy compared to questions in other subjects. A possible solution would be to utilize multimodal techniques, to produce models that can be evaluated with multimodal reasoning datasets (18; 34). Secondly, due to the stepwise attribute of SCDPO, it is not very effective on solution formats consisting of pure code. It only works on solutions consisting of natural language chain of thought or interleaved natural language and code. A method to properly enhance pure code solutions needs to be derived, which we leave for future work. Thirdly, as with all language models, our models can potentially generate hallucinations or produce misleading solutions, which can have a negative effect. This can be mitigated with methods such as verification, which we also leave for future work.

### v3

@@ 1 Introduction
Large language models (LLMs) have shown great potential in mathematical problem-solving. Recently, Direct Preference Optimization (DPO; (27)) has emerged as a popular choice for aligning LLMs with relative feedback to improve the quality of generated text. Prior works (8; 25; 43) have demonstrated that reinforcement learning algorithms and DPO can improve the mathematical reasoning abilities of LLMs, making the generated reasoning process more controllable. Different from other tasks that need human or AI feedback, the final answer to a mathematical problem serves as a reliable way to judge the quality of the model’s response, since a mathematical problem typically has a single correct answer. As a result, the responses producing the correct final answers are desirable and can serve as the preferred samples, while the ones reaching incorrect final answers are undesirable and can serve as the dispreferred samples.
@@ 2 Related Work
Aligning LLMs Using Relative Feedback. Methods that align LLMs with human or AI annotated preference data have been used to improve performance on a variety of downstream tasks such as translation (16), summarization (32; 56), and instruction-following (24; 28). Reinforcement learning from human (or AI) feedback (8; 4) first trains a reward model, then uses reinforcement learning algorithms such as REINFORCE (41), PPO (29), or variants (28) to maximize the reward. To simplify the pipeline, several direct alignment methods (27; 1; 54) have been proposed. Among them, DPO (27) and several of its variants (25; 10; 18; 1) offer a way to optimize the reward function without having to train an extra reward model, proving highly effective on various tasks (34; 50).
Recently, these preference alignment methods have also been applied to mathematical problem-solving tasks. DeepSeekMath (30) and R3 (42) uses RL to improve mathematical accuracy, while ChatGLM-Math (43) and Process Reward Synthesizing (15) uses DPO to improve model’s mathematical generation quality. Process supervision (17; 37) uses stepwise preference of mathematical solution in its RL finetuning, which is highly effective but needs costly fine-grained human annotation. Our work offers a way to create stepwise error annotations of preferred and dispreferred solution pairs, and uses the data to improve DPO’s performance on mathematical problem-solving tasks.
@@ 3 Step-Controlled DPO Pipeline
We experiment with two solution formats: code-integrated solution format (55), and chain-of-thought solution format (39). For the code-integrated solution format, $t_{i}$ in $(t_{0},\ldots,t_{m})$ alternates between a code reasoning step and a natural language reasoning step, while for the chain-of-thought solution format, all the steps are in natural language. We primarily use the code-integrated solution format, as previous works (36; 20; 12) show that it results in higher accuracy than chain-of-thought. We finetune a Mistral-7B model with 34K samples of code-integrated solutions from GSM8K and 47K from MATH to create the initial model. We also validate our method on the chain-of-thought format using the off-the-shell MetaMath-Mistral-7B model, as well as MathCoder-Mistral-7B, which we trained using the MathCodeInstruct dataset (36).
@@ 5.2 Scaling of Data Amount on 20B Model
Training Settings. We use InternLM2-20B (6) as the foundation model, as it has demonstrated high performance in previous works (20; 6), even surpassing larger models such as Mixtral-8x7B (14) and Llama2-70B (33) in some cases. In the SFT stage, we finetune the model with a learning rate of $1.0\times 10^{-5}$ for 3 epochs, with a context length of 2048 tokens. In DPO and SCDPO training, we use a learning rate of $1.5\times 10^{-7}$ to train the SFT model for 2 epochs, with a context length of 1024 and $\beta$ set to 0.1. The models are trained on 16 NVIDIA A800 80GB GPUs with a batch size of 64.
@@ 6 Limitations and Future Work
Our work contains the following limitations, and we leave them for future work. Firstly, our work is conducted on purely linguistic models, which struggle to solve mathematical problems requiring an understanding of images. For example, questions in the geometry subject of the MATH dataset exhibit lower accuracy compared to questions in other subjects. A possible solution would be to utilize multimodal techniques, to produce models that can be evaluated with multimodal reasoning datasets (19; 35). Secondly, due to the stepwise attribute of SCDPO, it is not very effective on solution formats consisting of pure code. It only works on solutions consisting of natural language chain of thought or interleaved natural language and code. A method to properly enhance pure code solutions needs to be derived, which we leave for future work. Thirdly, as with all language models, our models can potentially generate hallucinations or produce misleading solutions, which can have a negative effect. This can be mitigated with methods such as verification, which we also leave for future work.